# From Persona to Personalization: A Survey on Role-Playing Language Agents

### v1

@@ 5.3 Construction of Character RPLAs
This method includes pre-training and supervised fine-tuning. In pre-training, LLMs learn from large-scale web corpus which includes vast amounts of literary works and encyclopedia entries. This provides LLMs with knowledge of a wide range of established characters, such as Hermione Granger and Socrates, enabling LLMs to readily role-play these characters. Supervised fine-tuning for RPLAs is be adopted to tailor LLMs to role-play specific characters (Shao et al., 2023; Yu et al., 2024), or to develop foundation models with refined role-playing capabilities utilizing datasets of diversified characters and scenarios (Li et al., 2023a; Wang et al., 2023g).
@@
This method directly provides LLMs with character data in the context, leveraging the in-context learning capability of advanced LLMs. This serves as a simple yet effective methodology for RPLA construction, and is hence widely adopted by recent RPLAs (Wang et al., 2023g; Zhou et al., 2023a). However, character data is often voluminous, and interaction data between RPLAs and users is also continuously produced during the interaction process. This makes it impractical to include all data for a character RPLA within the context limits of LLMs. Consequently, long-term memory modules are being increasingly incorporated into RPLA frameworks to manage the vast amount of character RPLA data (Li et al., 2023a; Wang et al., 2023g; Xu et al., 2024b). These modules store most character knowledge and interaction data in a database, and retrieve necessary information in relevant scenarios.
@@ 5.4 Evaluation of Character RPLAs
They assess how well a foundation model is capable of the role-playing task, regardless of the characters it role-plays. This evaluation can be structured into the following levels:
@@
    Role-playing Engagement: Basically, the LLMs should actively participate in the role-playing scenario. They should produce responses in dialogue format and exhibit deep immersion, avoiding out-of-character utterance such as “As an AI model”). Additionally, the RPLAs are expected to exhibit stable and consistent personalities across different turns (Shao et al., 2023), sessions (Wang et al., 2024c) and even language (Huang et al., 2023b).
@@
    Anthropomorphic Capabilities: RPLAs are expected to acquire cognitive, emotional and social intelligence towards human levels. Relevant dimensions include conversation attractiveness (Zhou et al., 2023a; Tu et al., 2024), theory of mind (Kosinski, 2023; Mao et al., 2023), empathy (Sorin et al., 2023), emotional intelligence (Huang et al., 2023a), and goal-driven social skills (Zhou et al., 2024b; Wang et al., 2024a). These capabilities are practically important for RPLAs to effectively serve as emotional companions for humans.
Character Fidelities
They evaluate how a specific RPLA reproduces the intended character, which depends on both the foundation model, the agent framework, and the character data. Relevant dimensions are categorized into four categories: linguistic style and knowledge, which are considered superficial, as well as personality and thought, which represent deeper, underlying aspects:
@@
    Linguistic Style: Basically, RPLAs should speak in a tone that emulates the linguistic style of the intended characters (Wang et al., 2023g; Li et al., 2023a; Zhou et al., 2023a; Yu et al., 2024). For this purpose, RPLAs are typically provided with demonstrative character dialogues (Wang et al., 2023g; Li et al., 2023a), and they could mimic the tone leveraging the in-context learning ability of LLMs.
@@
    Knowledge: RPLAs are essentially required to simulate the character’s breadth of knowledge. On one hand, they should accurately recall knowledge of the character, including their identity (Zhou et al., 2023a; Wang et al., 2023g; Tang et al., 2024; Lu et al., 2024), social relationships (Chen et al., 2023c; Shen et al., 2023a; Zhao et al., 2023a), and experiences (Shao et al., 2023; Wang et al., 2023g; Chen et al., 2023c; Yu et al., 2024). On the other hand, they may be required to refrain from demonstrating knowledge or ability beyond the character’s scope (e.g., an LLM could write code even if it is role-playing Socrates, which is unnecessarily expected) (Shao et al., 2023; Lu et al., 2024; Yu et al., 2024). This phenomenon is referred to as “character hallucination” (Shao et al., 2023), which originates from the extensive knowledge possessed by LLMs and could be reduced via SFT (Shao et al., 2023).
@@
    Personality and Thinking Process: RPLAs are expected to capture the inner world of the characters, which can be measured upon their thoughts in concrete scenarios (Xu et al., 2024b; Chen et al., 2024) and their underlying personalities (Wang et al., 2024c; Shao et al., 2023). Advanced RPLAs should be able to understand and replicate how characters would think in specific scenarios, e.g., understanding their motivations for decisions (Yuan et al., 2024b), or predicting decisions and behaviors that align closely with the characters (Xu et al., 2024b; Chen et al., 2024). Personality is behind the concrete thoughts. It is the interrelated behavioral, cognitive and emotional patterns of individuals (Barrick & Mount, 1991; Bem, 1981), which applies to both characters and RPLAs. Hence, RPLAs should exhibit personality traits that match those of the characters (Wang et al., 2024c), which could be measured via psychological scales such as the Big Five Inventory.
@@
    Automatic Evaluation with Ground Truth: Typically, datasets with ground truth are expected for evaluating character fidelity in terms of knowledge, personality and thought. While early similarity metrics such as Rouge-L (Lin, 2004) could be applied to compare RPLA responses with ground truth (Wang et al., 2023g), recent studies increasingly leverage state-of-the-art LLMs such as GPT-4 as evaluators. On one hand, evaluator LLMs can score RPLA responses based on certain criteria, or determine the superior response from two models for win rate calculation (Wang et al., 2023g), provided with ground truth reference (Wang et al., 2023g). On the other hand, evaluator LLMs can be used to classify RPLA responses, and the results are then compared with ground truth labels (Wang et al., 2024c).
@@
    Automatic Evaluation without Ground Truth: As collecting ground truth data for RPLA evaluation is often challenging, several studies such as CharacterEval explore using LLMs to evaluate RPLA responses without ground truth (Shao et al., 2023; Tu et al., 2024). Instead, character profiles should be provided. This approach is effective for evaluating character-independent abilities and linguistic styles, which require little knowledge about the characters. However, when it comes to characters’ knowledge and thoughts, LLMs might not possess the necessary depth of relevant knowledge, especially for unfamiliar characters. This concern potentially leads to inadequately informed judgments of LLMs, and hence produces suboptimal evaluation results.
@@
    Multi-choice Questions: Multi-choice questions also come with ground truth, yet they differ from “automatic evaluation with ground truth” in that they merely require RPLAs to select from a fixed set of options, rather than generating open-ended responses. This significantly reduces the output space for RPLAs, making the evaluation simpler. This method is particularly suitable for evaluating the fidelity of characters’ thoughts, e.g., behavior prediction (Xu et al., 2024b; Chen et al., 2024) and motivation generation (Yuan et al., 2024b; Shen et al., 2023a). For these tasks, it is impractical to require RPLAs to produce responses exactly matching the ground truth, and responses may be reasonable even if they deviate from the ground truth significantly.
@@
    Human Evaluation: Inviting human annotators to assess the performance of RPLAs is a viable and effective approach (Zhou et al., 2023a). However, it comes with several drawbacks, such as cost in terms of time and money, as well as lack of reproducibility. This method is akin to “automatic evaluation without ground truth”, yet employs humans as more precise evaluators. Hence, it similarly falls short in evaluations that require in-depth knowledge about the characters, as recruiting qualified annotators who are well-acquainted with these characters can be difficult.
@@ 6.1 Definition
    Task Solving: Furthermore, personalized RPLAs become increasingly competent in more complicated task solving (Yao et al., 2023a; Significant-Gravitas, 2023), such as coding (Microsoft, 2024), travel planning (Xie et al., 2024a), and research survey (Wang et al., 2024b), typically interacting with various external software. They are autonomous LLM-based agents that are deeply integrated with personal data, devices, and services (Dong et al., 2023; Li et al., 2024d). They have significantly advanced personal assistants beyond early predecessors such as Siri (Apple Inc., 2024) which struggle with complex user requests.
@@ 6.2 Data Collection of Individualized Persona
Datasets #Profile #Interactions Domain Language Source
@@
Profiles are fundamental information that explicitly describes individualized personas, which are typically well-structured. Typically, they are initially set by users, and can be continuously updated. The basic elements usually include the names, gender and ethnicity of individual users in text (Santurkar et al., 2023) Besides, profiles commonly contain natural language descriptions of individuals, describing their characteristics, such as identity, hobbies, experiences and other statements (Zhang et al., 2018; Dinan et al., 2020; Gao et al., 2023b; Li et al., 2021; Ng et al., 2024), varying based on the detailed applications. For example, in live streaming applications, persona data can be composed of both basic profile information — such as an individual’s age, gender, and location — and domain-specific details, namely streamer characteristics such as fan numbers and streaming style Gao et al. (2023b). Additionally, profiles can contain multi-modal information. For instance, profiles in  (Ahn et al., 2023) incorporate text-image pairs, which are individuals’ comments for pictures on social media.
@@
The interaction data capture the dynamic evolution of individualized persona. Interactions are data generated during the use of applications that implicitly portray individualized personas, such as conversations, user preferences, and other behaviors. For example, PERSONA-CHAT (Zhang et al., 2018) and ConvAI (Dinan et al., 2020) collect two-person dialogues through crowd-sourcing, while LiveChat (Gao et al., 2023b) and MPCHAT (Ahn et al., 2023) collect multiplayer conversations from Internet sources such as live streaming and Reddit. To reduce the construction cost,  Jandaghi et al. (2023) adopts LLMs for dialogue synthesis. In addition to dialogues in natural language,  Agrawal et al. (2023) and  Santurkar et al. (2023) introduce comic pictures and multiple-choice questions as interactions. This kind of data could be consistently collected and systematically organized in real-world applications, offering benefits such as convenient acquisition and dynamic evolution. Hence, it plays an important role in practical applications.
@@ 6.3 Modeling Individualized Persona
This method tailor the outputs of LLMs to reflect specific personas represented in pre-existing datasets. Parameter fine-tuning emerges as the mainstream approach for offline learning, typically based on SFT and RLHF (Mondal et al., 2024; Zheng et al., 2023c; Li et al., 2024b; Jang et al., 2023). For example, Mondal et al. (2024) proposes a two-stage approach for personalizing LLMs with profile and interaction datasets. In addition, some recent studies propose techniques with nonparametric learning for LLMs personalization. For instance, Shea & Yu (2023) introduces an offline RL framework with a persona consistency critic and variance reduction, while Weng et al. (2024) integrates embedding control vectors within the model’s activation states, allowing dynamic output adjustment for diverse personality traits. These methods exhibit several deficiencies: 1) they face a fundamental trade-off between accuracy and efficiency; 2) they are heavily reliant on the quality of datasets; 3) more crucially, they struggle to adapt to dynamic changes in persona data, limiting their real-world applicability.
@@
In online learning, the personas are dynamic and continuously evolving, i.e., regularly updated with incoming data, the user interactions in real-world applications. This enables personalized RPLAs to quickly adapt and stay relevant to user needs and preferences. With LLMs, effective persona learning is typically nonparametric and training-free, which only involves effective management of memory and context (Dalvi Mishra et al., 2022; Kim et al., 2024; Baek et al., 2023; Zhou et al., 2024c). For this demand, retrieval modules become indispensable, especially for LLMs with limited context window (Mysore et al., 2023; Sun et al., 2024). Moreover, methodologies for effective online learning methods consider not only natural language interactions, but also non-linguistic feedback from users(Ma et al., 2023a). Besides nonparametric methods, fine-tuning with online interactive data is also widely applied to online persona learning, including both SFT with mini-batches from on-the-fly user stream data  (Qin et al., 2024) and RLHF with real-time user feedback (Ding et al., 2023b; Bai et al., 2022). Nevertheless, significant challenges arise in accurately recognizing and learning the sparse persona-specific features from the noisy interaction data. Besides, the personas of real users may change over time, which poses further challenges for their effective modeling and updating. Therefore, for nonparametric methods, the effectiveness heavily relies on the mechanisms of memory management and retrieval.

### v2

@@ 5.3 Construction of Character RPLAs
This method includes pre-training and supervised fine-tuning. In pre-training, LLMs learn from large-scale web corpus which includes vast amounts of literary works and encyclopedia entries. This provides LLMs with knowledge of a wide range of established characters, such as Hermione Granger and Socrates, enabling LLMs to readily role-play these characters. Supervised fine-tuning for RPLAs is be adopted to tailor LLMs to role-play specific characters (Shao et al., 2023; Yu et al., 2024), or to develop foundation models with refined role-playing capabilities utilizing datasets of diversified characters and scenarios (Li et al., 2023a; Wang et al., 2024a).
@@
This method directly provides LLMs with character data in the context, leveraging the in-context learning capability of advanced LLMs. This serves as a simple yet effective methodology for RPLA construction, and is hence widely adopted by recent RPLAs (Wang et al., 2024a; Zhou et al., 2023a). However, character data is often voluminous, and interaction data between RPLAs and users is also continuously produced during the interaction process. This makes it impractical to include all data for a character RPLA within the context limits of LLMs. Consequently, long-term memory modules are being increasingly incorporated into RPLA frameworks to manage the vast amount of character RPLA data (Li et al., 2023a; Wang et al., 2024a; Xu et al., 2024c). These modules store most character knowledge and interaction data in a database, and retrieve necessary information in relevant scenarios.
@@ 5.4 Evaluation of Character RPLAs
This line of work assesses how well a foundation model is capable of the role-playing task, regardless of the characters it role-plays. According to different levels of interaction capabilities, we have considered basic role-playing abilities and conversational skills, progressing to more in-depth anthropomorphic capabilities matched with humans. These have been categorized into the following three levels:
@@
    Role-playing Engagement: Basically, the LLMs should actively participate in the role-playing scenario. They should produce responses in dialogue format and exhibit deep immersion, avoiding out-of-character utterance such as “As an AI model”). Additionally, the RPLAs are expected to exhibit stable and consistent personalities across different turns (Shao et al., 2023), sessions (Wang et al., 2024d) and even language (Huang et al., 2023b).
@@
    Anthropomorphic Capabilities: RPLAs are expected to acquire cognitive, emotional and social intelligence towards human levels. Relevant dimensions include conversation attractiveness (Zhou et al., 2023a; Tu et al., 2024), theory of mind (Kosinski, 2023; Mao et al., 2023), empathy (Sorin et al., 2023), emotional intelligence (Huang et al., 2023a), and goal-driven social skills (Zhou et al., 2024b; Wang et al., 2024b). These capabilities are practically important for RPLAs to effectively serve as emotional companions for humans.
Character Fidelity
This line of work evaluates how a specific RPLA reproduces the intended character, which depends on both the foundation model, the agent framework, and the character data. Relevant dimensions are categorized into four categories: linguistic style and knowledge, which are considered superficial, as well as personality and thought, which represent deeper, underlying aspects:
@@
    Linguistic Style: Basically, RPLAs should speak in a tone that emulates the linguistic style of the intended characters (Wang et al., 2024a; Li et al., 2023a; Zhou et al., 2023a; Yu et al., 2024). For this purpose, RPLAs are typically provided with demonstrative character dialogues (Wang et al., 2024a; Li et al., 2023a), and they could mimic the tone leveraging the in-context learning ability of LLMs.
@@
    Knowledge: RPLAs are essentially required to simulate the character’s breadth of knowledge. On one hand, they should accurately recall knowledge of the character, including their identity (Zhou et al., 2023a; Wang et al., 2024a; Tang et al., 2024; Lu et al., 2024), social relationships (Chen et al., 2023c; Shen et al., 2023a; Zhao et al., 2023a), and experiences (Shao et al., 2023; Wang et al., 2024a; Chen et al., 2023c; Yu et al., 2024). On the other hand, they may be required to refrain from demonstrating knowledge or ability beyond the character’s scope (e.g., an LLM could write code even if it is role-playing Socrates, which is unnecessarily expected) (Shao et al., 2023; Lu et al., 2024; Yu et al., 2024). This phenomenon is referred to as “character hallucination” (Shao et al., 2023), which originates from the extensive knowledge possessed by LLMs and could be reduced via SFT (Shao et al., 2023).
@@
    Personality and Thinking Process: RPLAs are expected to capture the inner world of the characters, which can be measured upon their thoughts in concrete scenarios (Xu et al., 2024c; Chen et al., 2024) and their underlying personalities (Wang et al., 2024d; Shao et al., 2023). Advanced RPLAs should be able to understand and replicate how characters would think in specific scenarios, e.g., understanding their motivations for decisions (Yuan et al., 2024b), or predicting decisions and behaviors that align closely with the characters (Xu et al., 2024c; Chen et al., 2024). Personality is behind the concrete thoughts. It is the interrelated behavioral, cognitive and emotional patterns of individuals (Barrick & Mount, 1991; Bem, 1981), which applies to both characters and RPLAs. Hence, RPLAs should exhibit personality traits that match those of the characters (Wang et al., 2024d), which could be measured via psychological scales such as the Big Five Inventory.
@@
    Automatic Evaluation with Ground Truth: Typically, datasets with ground truth are expected for evaluating character fidelity in terms of knowledge, personality and thought. While early similarity metrics such as Rouge-L (Lin, 2004) could be applied to compare RPLA responses with ground truth (Wang et al., 2024a), recent studies increasingly leverage state-of-the-art LLMs such as GPT-4 as evaluators. On one hand, evaluator LLMs can score RPLA responses based on certain criteria, or determine the superior response from two models for win rate calculation, provided with “ground truth” responses as references. However, the “ground truth” are typically synthesized by advanced LLMs (Wang et al., 2024a) . On the other hand, evaluator LLMs can be used to classify RPLA responses, and the results are then compared with ground truth labels (Wang et al., 2024d).
@@
    Automatic Evaluation without Ground Truth: As collecting ground truth data for RPLA evaluation is often challenging, several studies such as CharacterEval explore using LLMs to evaluate RPLA responses without ground truth (Shao et al., 2023; Tu et al., 2024). Instead, character profiles should be provided. This category is effective for evaluating character-independent abilities and linguistic styles, which require little knowledge about the characters. However, when it comes to characters’ knowledge and thoughts, LLMs might not possess the necessary depth of relevant knowledge, especially for unfamiliar characters. This concern potentially leads to inadequately informed judgments of LLMs, and hence produces suboptimal evaluation results. Even when equipped with extensive knowledge, current LLMs still exhibit deficiencies in discerning nuanced aspects of role-playing.
@@
    Multi-choice Questions: Multi-choice questions also come with ground truth, yet they differ from “automatic evaluation with ground truth” in that they merely require RPLAs to select from a fixed set of options, rather than generating open-ended responses. This significantly reduces the output space for RPLAs, making the evaluation simpler. This method is particularly suitable for evaluating the fidelity of characters’ thoughts, e.g., behavior prediction (Xu et al., 2024c; Chen et al., 2024) and motivation generation (Yuan et al., 2024b; Shen et al., 2023a). For these tasks, it is impractical to require RPLAs to produce responses exactly matching the ground truth, and responses may be reasonable even if they deviate from the ground truth significantly.
@@
    Human Evaluation: Inviting human annotators to assess the performance of RPLAs is a viable and effective approach (Zhou et al., 2023a). However, it comes with several drawbacks, such as cost in terms of time and money, as well as lack of reproducibility. This method is akin to “automatic evaluation without ground truth”, yet employs humans as more precise evaluators. Hence, it similarly falls short in evaluations that require in-depth knowledge about the characters, as recruiting qualified annotators who are well-acquainted with these characters can be difficult. Combining automatic evaluation with human evaluation, some efforts also fine-tune the evaluator LLMs with human annotations (Tu et al., 2024).
Previous efforts primarily focus on RPLA evaluation on widely-known established characters (Wang et al., 2024d; Ahn et al., 2024). However, it is challenging to obtain high-quality datasets and achieve precise and nuanced evaluations for those highly complicated characters. Hence, there has recently been a growing trend in assessing LLMs’ role-playing capabilities based on original characters (BosonAI, 2024; Samuel et al., 2024).
@@ 6.1 Definition
    Task Solving: Furthermore, personalized RPLAs become increasingly competent in more complicated task solving (Yao et al., 2023a; Significant-Gravitas, 2023), such as coding (Microsoft, 2024), travel planning (Xie et al., 2024b), and research survey (Wang et al., 2024c), typically interacting with various external software. They are autonomous LLM-based agents that are deeply integrated with personal data, devices, and services (Dong et al., 2023; Li et al., 2024d). They have significantly advanced personal assistants beyond early predecessors such as Siri (Apple Inc., 2024) which struggle with complex user requests.
@@ 6.2 Data Collection of Individualized Persona
Datasets #Profile #Interactions Domain Lang. Source
@@
Profiles are fundamental information that explicitly describes individualized personas, which are typically well-structured. Typically, they are initially set by users, and can be continuously updated. The basic elements usually include the names, gender and ethnicity of individual users in text (Santurkar et al., 2023) Besides, profiles commonly contain natural language descriptions of individuals, describing their characteristics, such as identity, hobbies, experiences and other statements (Zhang et al., 2018; Dinan et al., 2020; Gao et al., 2023b; Li et al., 2021; Ng et al., 2024), varying based on the detailed applications. Additionally,  Lee et al. (2024) propose a multi-modality persona that includes elements such as the user’s appearance. For example, in live streaming applications, persona data can be composed of both basic profile information — such as an individual’s age, gender, and location — and domain-specific details, namely streamer characteristics such as fan numbers and streaming style Gao et al. (2023b). Additionally, profiles can contain multi-modal information. For instance, profiles in  (Ahn et al., 2023) incorporate text-image pairs, which are individuals’ comments for pictures on social media.
@@
The interaction data capture the dynamic evolution of individualized persona. Interactions are data generated during the use of applications that implicitly portray individualized personas, such as conversations, user preferences, and other behaviors. For example, PERSONA-CHAT (Zhang et al., 2018) and ConvAI (Dinan et al., 2020) collect two-person dialogues through crowd-sourcing, while LiveChat (Gao et al., 2023b) and MPCHAT (Ahn et al., 2023) collect multiplayer conversations from Internet sources such as live streaming and Reddit. To reduce construction costs,  Jandaghi et al. (2023) uses a Generator-Critic architecture with LLMs for dialogue synthesis, and  Zeng et al. (2024) first conducts automatic annotation to generate conversational data from raw sources like experiences, speeches, or writings. In addition to dialogues in natural language,  Agrawal et al. (2023) and  Santurkar et al. (2023) introduce comic pictures and multiple-choice questions as interactions. This kind of data could be consistently collected and systematically organized in real-world applications, offering benefits such as convenient acquisition and dynamic evolution. Hence, it plays an important role in practical applications.
@@ 6.3 Modeling Individualized Persona
This method tailors the outputs of LLMs to reflect specific personas represented in pre-existing datasets. Parameter fine-tuning emerges as the mainstream approach for offline learning, typically based on SFT and RLHF (Mondal et al., 2024; Zheng et al., 2023c; Li et al., 2024b; Jang et al., 2023). For example, Mondal et al. (2024) proposes a two-stage approach for personalizing LLMs with profile and interaction datasets. In addition, some recent studies propose techniques with nonparametric learning for LLMs personalization. For instance, Shea & Yu (2023) introduces an offline RL framework with a persona consistency critic and variance reduction, while Weng et al. (2024) integrates embedding control vectors within the model’s activation states, allowing dynamic output adjustment for diverse personality traits. These methods exhibit several deficiencies: 1) they face a fundamental trade-off between accuracy and efficiency; 2) they are heavily reliant on the quality of datasets; 3) more crucially, they struggle to adapt to dynamic changes in persona data, limiting their real-world applicability.
@@
In online learning, the personas are dynamic and continuously evolving, i.e., regularly updated with incoming data, the user interactions in real-world applications. This enables personalized RPLAs to quickly adapt and stay relevant to user needs and preferences. With LLMs, effective persona learning is typically nonparametric and training-free, which only involves effective management of memory and context (Dalvi Mishra et al., 2022; Kim et al., 2024; Baek et al., 2023; Zhou et al., 2024c). For this demand, retrieval modules become indispensable, especially for LLMs with limited context window (Mysore et al., 2023; Sun et al., 2024). Moreover, methodologies for effective online learning methods consider not only natural language interactions, but also non-linguistic feedback from users (Ma et al., 2023a).
Besides nonparametric methods, fine-tuning with online interactive data is also widely applied to online persona learning, including both SFT with mini-batches from on-the-fly user stream data  (Qin et al., 2024) and RLHF with real-time user feedback (Ding et al., 2023b; Bai et al., 2022a). Additionally,  Shaikh et al. (2024b) use fewer than 10 demonstrations to align language model outputs with a user’s demonstrated behaviors through iterative DPO (Rafailov et al., 2024) training. Nevertheless, significant challenges arise in accurately recognizing and learning the sparse persona-specific features from the noisy interaction data. Besides, the personas of real users may change over time, which poses further challenges for their effective modeling and updating. Therefore, for nonparametric methods, the effectiveness heavily relies on the mechanisms of memory management and retrieval.